# Flash Inference: Near Linear Time Inference for Long Convolution Sequence Models and Beyond

### v1

@@ Abstract
While transformers have been at the core of most recent advancements in sequence generative models, their computational cost remains quadratic in sequence length. Several subquadratic architectures have been proposed to address this computational issue. Some of them, including long convolution sequence models (LCSMs), such as Hyena, address this issue at training time but remain quadratic during inference. We propose a method for speeding up LCSMs’ exact inference to quasilinear $O(L\log^{2}L)$ time, identify the key properties that make this possible, and propose a general framework that exploits these. Our approach, inspired by previous work on relaxed polynomial interpolation, is based on a tiling which helps decrease memory movement and share computation. It has the added benefit of allowing for almost complete parallelization across layers of the position-mixing part of the architecture. Empirically, we provide a proof of concept implementation for Hyena, which gets up to 1.6×1.6$\times$1.6 × end-to-end improvement over standard inference by improving 50×50$\times$50 × within the position-mixing part.
@@ 1 Introduction
A lot of recent progress in deep learning, particularly in the form of large language models (LLMs) has been driven by the transformer architecture (Vaswani et al., 2017). While these models have great quality, it comes at a computation cost which scales quadratically in sequence length - both during training and inference. This can become prohibitive for very long contexts and as such a number of alternative architectures with better computational scaling in context length have been proposed (Gu and Dao, 2023; Poli et al., 2023; Fu et al., 2024). While most of these works have improved computational efficiency for training, some still scale quadratically in sequence length when it comes to inference, thus not improving asymptotically over transformers.
In this work, we propose a framework for optimizing inference efficiency for a general class of such models. As a case study, which inspired the method, we focus on long convolution sequence models (LCSMs) (Poli et al., 2023; Fu et al., 2022; Romero et al., 2021; Li et al., 2022; Karami and Ghodsi, 2024; Fu et al., 2023a). However, our approach is not limited to LCSMs alone and we identify the properties that allow for such inference speedups in hope to guide the design of future architectures.
In the particular case of LCSMs (including Hyena), the building block of the architecture is that of convolving the input sequence with a sequence-length long, (potentially underparameterized) filter. If we let $L$ be the sequence length (e.g. number of tokens in the case of LLMs), then a naive implementation of convolution during training would take $\Omega(L^{2})$ FLOPs, but one can employ FFT to bring that down to $O(L\log L)$. The issue that occurs during inference is that FFT cannot be used directly since the whole input sequence is not known ahead of time, but rather incrementally computed. Because of this, the naive inference approach goes up to $\Omega(L^{2})$ this is the apparent cost of moving from a static input to a dynamic one.
@@
  • Proposing the first quasilinear-in-sequence-length $O(L\log^{2}L)$, exact inference algorithm for LCSMs.
@@
  • Saving up on data movement thanks to the tiling approach used. We show how to save factors of $2$ in several places, including activation storage, when the convolutional filters are data-independent. However, our method extends to any causal, data-dependent filters.
  • Our framework provides an $O(L\log^{2}L)$ inference algorithm which, when run empirically, yields up to 1.6×1.6$\times$1.6 × end-to-end time-efficiency improvement; and up to 50×50$\times$50 × on the long convolution part.
Figure 1: Cell ($i$, $j$) corresponds to the contribution of mixer-input $y_{i}$ to mixer-output $z_{j}$. To compute $z_{j}$, all its line of contributions should be be accounted for. Because of the autoregressive nature of inference, one only has access to $y_{i}$ after $z_{i-1}$ has been computed. (Left Top) represents the standard (lazy) approach, (Left Bottom) represents the eager approach, and (Right) represents our suggested method.
@@ 2 Setting and Related Work
Much recent effort has been put in designing more efficient alternatives to transformers (Tay et al., 2022). A couple of well-performing classes are state space models (SSMs) (Gu et al., 2021; Gu and Dao, 2023) and convolution-based models(Fu et al., 2022; Poli et al., 2023). In this work we focus on the latter class. To avoid a popular misconception, it is worth emphasizing that while any linear-time invariant (LTI) SSM has an equivalent convolutional filter, it is not true that any convolutional filter has an equivalent low-dimensional LTI SSM representation. It could be the case that the smallest LTI SSM equivalent to a length-$L$ convolution filter has a dimension of $L$. However, such a dimension would deem an LTI SSM approach have quadratic time complexity both during training as well as during inference.
@@ 2.1 Notations and Definitions
In the most general form, we consider architectures obtained by stacking several position-mixing layers potentially interleaved with element-wise, feature-mixing modules, partially following Arora et al. (2023)’s notation. Put formally, let $\mathrm{block}_{1},\ldots\mathrm{block}_{M}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ be feature-mixing modules and $\mathrm{mixer}_{1}\ldots\mathrm{mixer}_{M}:\mathbb{R}^{L\times D}\to\mathbb{R}^{L\times D}$ be position-mixing modules. One can think of the blocks as being combinations of MLPs, layernorms, skip connections and other modules of similar purpose and of mixers as playing the role of multi-head attention (MHA) in a transformer - the part where embeddings at different positions interact with each other. Assume these are all learnt from some well-defined parameterized classes. Here, $M$ is the number of layers, $D$ is the used embedding dimension and $L$ can be any sequence length (unless positional embeddings are used, case in which it can only take values up to some $L_{max}$). We define the activations $a_{0},\ldots,a_{M}:\mathbb{R}^{L\times D}\to\mathbb{R}^{L\times D}$ at level $1\leq\ell\leq M$ and position $1\leq i\leq L$ as $(a_{\ell}(x))_{i}\triangleq\mathrm{block}_{\ell}(\mathrm{mixer}_{\ell}(a_{\ell-1}(x))_{i})$ where $a_{0}(x)=x$ represents the input embeddings ($L$ $D$-dimensional tokens) and the model output at position $i$ is read from the last layer $(a_{M}(x))_{i}$.
We focus on autoregressive sequence models where for an input sequence $x_{1},\ldots x_{p}$, one generates the $(p+1)^{\text{th}}$ token by sampling from a distribution given by $(a_{M}(x))_{p}$ (usually via a linear map). In order for this generation procedure to be sensible, we constrain (by construction of $\mathrm{mixer}$s) the models to be causal, that is, $(a_{\ell}(x))_{i}$ is a function of only $x_{[1,i]}$ or, equivalently, $\mathrm{mixer}_{\ell}(y)_{i}$ is a function of $y_{[1..i]}$.
We use $\odot$ to denote the element-wise (Hadamard) product: for two matrices $A,B\in\mathbb{R}^{N\times M}$, $A\odot B\in\mathbb{R}^{N\times M}$ is given by $(A\odot B)_{i,j}=A_{i,j}\cdot B_{i,j}$. We also use indexing by ranges to mean slicing a tensor in the respective dimension according to the range and indexing by a dot to mean leaving the dimension unchanged: for example, if $A:\mathbb{R}^{N\times M\times K}$, then $A_{\cdot,[l,r],5}\in\mathbb{R}^{N\times[r-l+1]}$ and $(A_{\cdot,[l,r],5})_{i,j}=A_{i,j+l-1,5}$.
We call a function $f:{\mathcal{X}}^{*}\to{\mathcal{X}}$ associative, if $f(x_{1},x_{2},\ldots x_{p})=f(f(x_{1},\ldots x_{i}),f(x_{i+1},\ldots x_{p}))$ for any $1\leq i<p$ and $x_{1}\ldots x_{p}\in{\mathcal{X}}$ that is, it is invariant to bracketing. Finally, we define the time complexity of an algorithm as the number of floating-point operations (FLOPs) it executes - this is a notion independent of hardware.
@@ 2.2 Self-attention
Self-attention is the building block of transformers (Vaswani et al., 2017). In its simplest form, it implements a $\mathrm{mixer}$ by using three projection matrices $Q,K,V\in\mathbb{R}^{D\times D}$ to obtain three sequences of vectors $q,k,v\in\mathbb{R}^{L\times D}$ by projecting the input $y\in\mathbb{R}^{L\times D}$ via $q=yQ,k=yK,v=yV$ these are called queries, keys and values, respectively. The (causal) self-attention operator $attention:\mathbb{R}^{L\times D}\to\mathbb{R}^{L\times D}$ is then given by:
@@ 2.3 Long Convolution Sequence Models (LCSMs)
LCSMs (Poli et al., 2023; Li et al., 2022; Gaido et al., 2024) work by creating a SISO (single-input-single-output) primitive to map an input sequence $y\in\mathbb{R}^{L}$ to $z\in\mathbb{R}^{L}$ via $z_{t}\mapsto\sum_{i=1}^{t}{y_{i}\cdot\rho_{t-i}}$ where $\rho$ is a (possibly infinite, and of length at least $L$) convolution filter which is often parameterized by a smaller latent space $\Theta$: $\rho=f(\theta)$ where $\theta\in\Theta,\dim{\Theta}\ll L$ is learnt. These convolution primitives operate on independent axes of the hidden dimensions to create a positional mixer: $\mathrm{mixer}(y)_{t,c}=\sum_{i=1}^{t}{y_{i,c}\cdot\rho_{t-i,c}}$. This assumes the filters to be independent, but shared ones are possible as well (as is the case for multi-head Hyena (Massaroli et al., 2024)).
@@ 2.3.1 The Inference Problem
Whereas during training, the convolutions can be performed in $O(L\log L)$ time via FFT as the whole of $y$ is known beforehand, during inference this is non-trivial. Suppose there is a prompt of size $P$ and we want to generate $L-P$ more tokens. The bottleneck problem can be understood as computing, for every layer and every dimension, the below:
@@
for all $1\leq t\leq L$ To pre-fill the first $P$ values $z_{[1..p]}$ at all levels of activations, since the first $P$ inputs are known, one can perform FFT as they would normally do during train-time (incurring an $O(P\log P)$ time cost), but past $P$th position it is important to note that $y_{i}$ is not available before computing $z_{i-1}$ this is where static FFT breaks. Since dealing with the prompt can be done easily (Massaroli et al., 2024, Lemma 2.1) by essentially filling in all contributions of $y_{[1..P]}$ to $z_{[1..L]}$ and then forgetting the prompt ever existed, we henceforth assume $P=0$.
@@ 2.3.2 Previous Work on Efficient Inference
The downside, however, is that by definition this will only be an approximation of the learnt filter. More importantly, this approximation represents a projection to a fundamentally smaller space of models, thus defeating the purpose of using LCSMs instead of LTI SSMs - it is practically a different training procedure for an LTI SSM model. Furthermore, this approach assumes the filter $\rho$ is data-independent - this is needed in order to undergo an expensive distillation procedure as a precomputation step. This can be qualitatively limiting as shown in Arora et al. (2023). While we do store all activations, our proposed framework exactly simulates the architecture and data-dependent filters can also be accommodated.
@@ 3 Fast LCSM inference
In this section we describe our method for the case of long convolution sequence models. That is, we still work with models as described in Section 2.1, but we further assume the mixers to be convolution-based, as described in Section 2.3. Hence, for every $\mathrm{mixer}_{\ell}\in\{\mathrm{mixer},\ldots\mathrm{mixer}_{M}\}$, there exists a filter $\rho_{\ell}\in R^{L\times D}$ such that
$\mathrm{mixer}_{\ell}(y)_{t}=\sum_{i=1}^{t}{y_{i}\odot\rho_{\ell,t-i}}$
We assume here, for simplicity, that $\rho$ is data-independent and part of the model, as is the case for Hyena. However, we discuss in Appendix B how our method can be extended to data-dependent filters. We also assume $L=2^{P}$ for some integer $P$ for simplicity - one can always round $L$ up to the closest power of $2$ without changing it by more than a factor of $2$ and thus keeping the asymptotics of all our analysis unchanged.
@@ 3.1.1 Relaxed Polynomial Interpolation
The problem of relaxed polynomial interpolation can be thought of as having to compute $z_{t}$ for every $1\leq t\leq L$ with the further constraint that $y_{t}$ and $\rho_{t}$ are only made available once $z_{t-1}$ has been outputted. While this general setting is covered in Appendix B, here we focus on the case that all of $\rho$ is known ahead of time and only $y$ is incrementally revealed. One could simply use the more general approach of (van der Hoeven, 1997) even when $\rho$ is known ahead of time, but we can take advantage of this particularity to get a twice-faster and slightly simpler algorithm. We present this both for brevity and because existing architectures have data-independent $\rho$’s so this is the more efficient choice for them (and thus, for practical purposes).
@@
    Hence, the eager approach can be thought of as performing work as soon as it can be performed. Here, $t$th iteration takes $O(L-t)$ so we still have an overall complexity of $O(L^{2})$.
@@
The employment of FFT for speeding up accounting for groups of contributions is formalized in the following lemma:
@@
Let $1\leq l\leq r\leq l^{\prime}\leq r^{\prime}\leq L$ represent ranges of lengths $L_{1}=r-l+1$ and $L_{2}=r^{\prime}-l^{\prime}+1$ of $y$ and $z$, respectively. There exists an FFT-based algorithm running in $O(L_{1}+L_{2})$ space and $O((L_{1}+L_{2})\log(L_{1}+L_{2}))$ time complexity that, given access to $y_{[l,r]}$, computes the aggregated contributions of all $y_{[l,r]}$ to all of $z_{[l^{\prime},r^{\prime}]}$. That is, for each $l^{\prime}\leq t\leq r^{\prime}$, it computes:
for all $l^{\prime}\leq t\leq r^{\prime}$.
@@
The lazy, eager, and our version of a relaxed approach are depicted in Figure 1 under the form of tilings of the contribution space. Note that the lazy and eager approaches can still be thought of as grouping contribution accounting, but do so along thin strips: the contribution of a long range of $y$’s to one element of $z$ in the lazy approach and that of a singe element of $y$ to a long range of $z$’s in the eager approach. The key to our relaxed tiling is using more balanced tiles since, as noted, the cost of a tile is given by the larger side (which is very expensive for thin tiles). The tiles still have to respect $r\leq l^{\prime}$ and one needs to ensure that the whole line of contribution to some $z_{t}$ has been dealt with by the time we return the value of $z_{t}$ this is indeed the case for our tiling.
@@
2:  for $i\leftarrow 1$ to $L-1$ do
3:     $U\leftarrow$ maximum power of $2$ that divides $i$ the side of the $i$th gray tile
4:     $z_{i}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=y_{i}*\rho_{0}$ finalize $z_{i}$ by accounting for newly available $y_{i}$ red cell
5:     # account for the contribution of $y_{[i-U+1,i]}$ to $z_{[i+1,i+U]}$ gray tile
6:     $z_{[i+1,i+U]}\mathrel{+}=\tau(y,[i-U+1,i],\rho,[i+1,i+U])$ $\tau$ is defined as in Eq 7 and computed as per Lemma 1
7:     return $z_{i}$ $z_{i}$ has been computed
8:     $unlock(y_{i+1})$ $y_{i+1}$ becomes available
@@
10:  $z_{L}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=y_{L}*\rho_{0}$
11:  return $z_{L}$ $z_{L}$ has been computed
While the lazy and eager algorithms have been described, one needs to make explicit the order of processing the tiles in the relaxed setting since we still cannot use $y_{t}$ before $z_{t-1}$ has been computed. This is shown in Algorithm 1: return marks a value of $z$ having been computed and unlock represents a value of $y$ becoming available. At the beginning of iteration $i$ it always holds that $z_{i}=\sum_{j=1}^{i-1}{y_{j}\cdot\rho_{i-j}}$. We start the iteration by completing this with the freshly unlocked $y_{i}\cdot\rho_{0}$ at line 4 - this corresponds to the red tiles of Figure 1. Following this step $z_{i}$ is ready to be returned. Then, we account for the gray tile that has just been unlocked (line 6) - this tile corresponds to the contribution of $y_{[i-U+1,i]}$ to $z_{[i+1,i+U]}$ where $U$ is the side of the tile, namely the largest power of $2$ dividing $i$. Finally we return $z_{i}$ as being in its final form and get access to $y_{i+1}$ we could have done this before dealing with the gray tile as well. Note that the relative ordering of the red and gray lines does not matter within an iteration, but it is important to finish up the first red and gray tiles before second ones, the second ones before the third ones and so on.

### v2

@@ Abstract
While transformers have been at the core of most recent advancements in sequence generative models, their computational cost remains quadratic in sequence length. Several subquadratic architectures have been proposed to address this computational issue. Some of them, including long convolution sequence models (LCSMs), address this issue at training time but remain quadratic during inference. We propose a method for speeding up LCSMs’ exact inference to quasilinear time, identify the key properties that make this possible, and propose a general framework that exploits these. Our approach, inspired by previous work on relaxed polynomial interpolation, is based on a tiling which helps decrease memory movement and share computation. It has the added benefit of allowing for almost complete parallelization across layers of the position-mixing part of the architecture. Empirically, we provide a proof of concept implementation for Hyena, which gets up to $7.8\times$ end-to-end improvement over standard inference by improving up to $110\times$ within the position-mixing part.
@@ 1 Introduction
A lot of recent progress in deep learning, particularly in the form of large language models (LLMs) has been driven by the transformer architecture (Vaswani et al., 2017). While these models have great quality, it comes at a computation cost which scales quadratically in sequence length - both during training and inference. This can become prohibitive for very long contexts and as such a number of alternative architectures with better computational scaling in context length have been proposed (Gu and Dao, 2023; Poli et al., 2023; Fu et al., 2024). Although most of these works have improved computational efficiency for training, some still scale quadratically in sequence length when it comes to inference, thus not improving asymptotically over transformers.
In this work, we propose a framework for optimizing inference efficiency for a general class of such models. As a case study, which inspired the method, we focus on long convolution sequence models (LCSMs) (Poli et al., 2023; Fu et al., 2022; Romero et al., 2021b; Li et al., 2022; Karami and Ghodsi, 2024; Fu et al., 2023a; Romero et al., 2021a). However, our approach is not limited to LCSMs alone and we identify the properties that allow for such inference speedups in hope to guide the design of future architectures.
In the particular case of LCSMs, the building block of the architecture is that of convolving the input sequence with a sequence-length long, (potentially underparameterized) filter. If we let $L$ be the sequence length (e.g. number of tokens in the case of LLMs), then a naive implementation of convolution during training would take $\Omega(L^{2})$ FLOPs, but one can employ FFT to bring that down to $O(L\log L)$. The issue that occurs during inference is that FFT cannot be used directly since the whole input sequence is not known ahead of time, but rather incrementally computed. Because of this, the naive inference approach goes up to $\Omega(L^{2})$ - this is the apparent cost of moving from a static input to a dynamic one.
@@
Figure 1: Cell (ii, $j$) corresponds to the contribution of mixer-input $y_{i}$ to mixer-output $z_{j}$. To compute $z_{j}$, all its line of contributions should be be accounted for. Because of the autoregressive nature of inference, one only has access to $y_{i}$ after $z_{i-1}$ has been computed. (Left Top) represents the standard (lazy) approach, (Left Bottom) represents the eager approach, and (Right) represents our suggested method.
@@
  • Proposing the first quasilinear-in-sequence-length $O(L\log^{2}L)$, exact inference algorithm for LCSMs. This is in contrast to previous work of (Massaroli et al., 2024) who performs approximate inference.
@@
  • Saving up on data movement thanks to the tiling approach used. We show how to save factors of $2$ in several places, including activation storage, when the convolutional filters are data independent. However, our method extends to any causal, data-dependent filters.
  • Our framework provides an $O(L\log^{2}L)$ inference algorithm, which, when run empirically, yields up to $7.8\times$ end-to-end time-efficiency improvement; and up to $110\times$ on the long convolution part.
@@ 2 Setting and Related Work
Recent efforts have focused on designing more efficient alternatives to transformers (Tay et al., 2022), with state space models (SSMs) (Gu et al., 2021a; Gu and Dao, 2023) and convolution-based models (Fu et al., 2022; Poli et al., 2023) emerging as notable contenders. In this work, we focus on the latter simply because we leverage the convoluional view directly. Nevertheless, it is worth mentioning that there is a one-to-one mapping between linear-time invariant (LTI) SSMs and LCSMs and consequently, our findings apply to LTI SSMs as well (since they represent the same models). While LCSMs inherently assume sequence-long convolutions, they can correspond to LTI SSMs of varying state dimensions (up to filter length). Using the recurrent formulation of SSMs for inference incurs costs that scale with the state dimension, which can become prohibitive for larger dimensions. In contrast, our method exclusively leverages the convolutional perspective and has no dependency on filter properties, achieving consistent computational performance regardless of the filter (or its equivalent state dimension). Thus, while the recurrent view may be more efficient for low-dimensional settings, our approach remains broadly applicable and computationally efficient for LTI SSMs of any dimension. An extended contextual discussion is available in Appendix A.
@@ 2.1 Notations and Definitions
In the most general form, we consider architectures obtained by stacking several position-mixing layers potentially interleaved with element-wise, feature-mixing modules, partially following Arora et al. (2023)’s notation. Put formally, let $\mathrm{block}^{1},\ldots\mathrm{block}^{M}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ be feature-mixing modules and $\mathrm{mixer}^{1}\ldots\mathrm{mixer}^{M}:\mathbb{R}^{L\times D}\to\mathbb{R}^{L\times D}$ be position-mixing modules (we use superscripts throughout to denote layers). One can think of the blocks as being combinations of MLPs, layernorms, skip connections and other modules of similar purpose and of mixers as playing the role of multi-head attention (MHA) in a transformer - the part where embeddings at different positions interact with each other. Assume that these are all learned from some well-defined parameterized classes. Here, $M$ is the number of layers, $D$ is the embedding dimension used, and $L$ can be any sequence length (unless positional embeddings are used, case in which it can only take values up to some $L_{max}$). We define the activations $a^{0},\ldots,a^{M}:\mathbb{R}^{L\times D}\to\mathbb{R}^{L\times D}$ at level $1\leq\ell\leq M$ and position $1\leq i\leq L$ as $a^{\ell}(x)_{i}\triangleq\mathrm{block}^{\ell}(\mathrm{mixer}^{\ell}(a^{\ell-1}(x))_{i})$ where $a^{0}(x)=x$ represents the input embeddings ($L$ $D$-dimensional tokens) and the model output at position $i$ is read from the last layer $a^{M}(x)_{i}$.
We focus on autoregressive sequence models where for an input sequence $x_{1},\ldots x_{p}$, one generates the $(p+1)^{\text{th}}$ token by sampling from a distribution given by $a^{M}(x)_{p}$ (usually via a linear map). In order for this generation procedure to be sensible, we constrain (by construction of $\mathrm{mixer}$s) the models to be causal, that is, $a^{\ell}(x)_{i}$ is a function of only $x_{[1,i]}$ or, equivalently, $\mathrm{mixer}^{\ell}(y)_{i}$ is a function of $y_{[1..i]}$.
We use $\odot$ to denote the element-wise (Hadamard) product: for two matrices $A,B\in\mathbb{R}^{N\times M}$, $A\odot B\in\mathbb{R}^{N\times M}$ is given by $(A\odot B)_{i,j}=A_{i,j}\cdot B_{i,j}$. We call a function $f:{\mathcal{X}}^{*}\to{\mathcal{X}}$ associative, if $f(x_{1},x_{2},\ldots x_{p})=f(f(x_{1},\ldots x_{i}),f(x_{i+1},\ldots x_{p}))$ for any $1\leq i<p$ and $x_{1}\ldots x_{p}\in{\mathcal{X}}$. Finally, we define the time complexity of an algorithm as the number of floating-point operations (FLOPs) it executes - this is a notion independent of hardware.
@@ 2.2 Self-attention
Self-attention is the building block of transformers (Vaswani et al., 2017). In its simplest form, it implements a $\mathrm{mixer}$ by using three projection matrices $Q,K,V\in\mathbb{R}^{D\times D}$ to obtain three sequences of vectors $q,k,v\in\mathbb{R}^{L\times D}$ by projecting the input $y\in\mathbb{R}^{L\times D}$ via $q=yQ,k=yK,v=yV$ - these are called queries, keys and values, respectively. The (causal) self-attention operator $attention:\mathbb{R}^{L\times D}\to\mathbb{R}^{L\times D}$ is then given by:
@@ 2.3 Long Convolution Sequence Models (LCSMs)
LCSMs (Poli et al., 2023; Li et al., 2022; Gaido et al., 2024) work by creating a SISO (single-input-single-output) primitive to map an input sequence $y\in\mathbb{R}^{L}$ to $z\in\mathbb{R}^{L}$ via $z_{t}\triangleq\sum_{i=1}^{t}{y_{i}\cdot\rho_{t-i}}$ where $\rho$ is a (possibly infinite, and of length at least $L$) convolution filter which is often parameterized by a smaller latent space $\Theta$: $\rho=f(\theta)$ where $\theta\in\Theta,\dim{\Theta}\ll L$ is learned. These convolution primitives operate on independent axes of the hidden dimensions to create a positional mixer: $\mathrm{mixer}(y)_{t}\triangleq\sum_{i=1}^{t}{y_{i}\odot\rho_{t-i}}\in\mathbb{R}^{D}$. This assumes the filters to be independent, but shared ones are possible as well (as is the case for multi-head Hyena (Massaroli et al., 2024)).
@@ 2.3.1 The Inference Problem
Whereas during training, the convolutions can be performed in $O(L\log L)$ time via FFT as the whole of $y$ is known beforehand, during inference this is non-trivial. Suppose there is a prompt of size $P$ and we want to generate LPL-P more tokens. The bottleneck problem can be understood as computing, for every layer and every dimension, the below:
@@
for all $1\leq t\leq L$ . To pre-fill the first $P$ values $z_{[1..p]}$ at all levels of activations, since the first $P$ inputs are known, one can perform FFT as they would normally do during train-time (incurring an $O(P\log P)$ time cost), but past $P$th position it is important to note that $y_{i}$ is not available before computing $z_{i-1}$ - this is where static FFT breaks. Since dealing with the prompt can be done easily (Massaroli et al., 2024, Lemma 2.1) by essentially filling in all contributions of $y_{[1..P]}$ to $z_{[1..L]}$ and then forgetting the prompt ever existed, we henceforth assume $P=0$.
@@ 2.3.2 Previous Work on Efficient Inference
The downside, however, is that by definition this will only be an approximation of the learned filter. More importantly, this approximation represents a projection to a potentially different space of models, making the end-to-end procedure practically a different training procedure for a low-dimensional LTI SSM model. Furthermore, this approach assumes the filter $\rho$ is data-independent - this is needed in order to undergo an expensive distillation procedure as a precomputation step. This can be qualitatively limiting as shown in Arora et al. (2023). While we do store all activations, our proposed framework exactly simulates the architecture and data-dependent filters can also be accommodated.
Finally, we note that concurrent work from Agarwal et al. (2024) approaches the same problem, with an emphasis on applications to linear dynamical systems, via a similar algorithm. Our frameworks goes farther by generalizing to data-dependent convolutional filters, while also tackling a number of systems-levels challenges in making the method practically relevant.
@@ 3 Fast LCSM inference
In this section we describe our method for the case of long convolution sequence models. That is, we still work with models as described in Section 2.1, but we further assume the mixers to be convolution-based, as described in 2.3. Hence, for every $\mathrm{mixer}^{\ell}\in\{\mathrm{mixer},\ldots\mathrm{mixer}^{M}\}$, there exists a filter $\rho^{\ell}\in R^{L\times D}$ such that
$\mathrm{mixer}^{\ell}(y)_{t}=\sum_{i=1}^{t}{y_{i}\odot\rho^{\ell}_{t-i}}$
We assume here, for simplicity, that $\rho$ is data-independent and part of the model, as is the case for most existing architectures. However, we discuss in Appendix B how our method can be extended to data-dependent filters (by paying a factor of $2$). We also assume $L=2^{P}$ for some integer $P$ for simplicity - one can always round $L$ up to the closest power of $2$ without changing it by more than a factor of $2$ and thus keeping the asymptotics of all our analysis unchanged.
@@ 3.1.1 Relaxed Polynomial Interpolation
The problem of relaxed polynomial interpolation can be thought of as having to compute $z_{t}$ for every $1\leq t\leq L$ with the further constraint that $y_{t}$ and $\rho_{t}$ are only made available once $z_{t-1}$ has been outputted. While this general setting is covered in Appendix B, here we focus on the case that all of $\rho$ is known ahead of time and only $y$ is incrementally revealed. Whereas one could simply use the more general approach of (van der Hoeven, 1997), we can further take advantage of $\rho$ being known ahead of time to get a twice-faster and slightly simpler algorithm. We present this both for brevity and because existing architectures have data-independent $\rho$’s so this is the more efficient choice for them (and thus, for practical purposes). However, (Arora et al., 2023; Karami and Ghodsi, 2024) show that data-dependent filters can improve quality but finding a good causal architecture remains open.
@@
    Hence, the eager approach can be thought of as performing work as soon as it can be performed. Here, $t$th iteration takes O(Lt)O(L-t) so we still have an overall complexity of $O(L^{2})$.
@@
First, let us define the sort of contribution grouping that we consider:
Definition.
For any $1\leq l\leq r\leq l^{\prime}\leq r^{\prime}\leq L$ we let $\tau(y,[l,r],\rho,[l^{\prime},r^{\prime}])$ represent the aggregated contribution of all $y_{[l,r]}$ to all of $z_{[l^{\prime},r^{\prime}]}$. That is, for all $l^{\prime}\leq t\leq r^{\prime}$, we have:
The employment of FFT for speeding up the computation of $\tau(y,[l,r],\rho,[l^{\prime},r^{\prime}])$ can then be achieved following the next lemma:
@@
Let $1\leq l\leq r\leq l^{\prime}\leq r^{\prime}\leq L$ represent ranges of lengths $L_{1}=r-l+1$ and $L_{2}=r^{\prime}-l^{\prime}+1$ of $y$ and $z$, respectively. There exists an FFT-based algorithm running in $O(L_{1}+L_{2})$ space and $O((L_{1}+L_{2})\log(L_{1}+L_{2}))$ time complexity that, given access to $y_{[l,r]}$, computes all the elements of $\tau(y,[l,r],\rho,[l^{\prime},r^{\prime}])_{[l^{\prime},r^{\prime}]}$ - that is, all the aggregated contributions of $y_{[l,r]}$ to all of $z_{[l^{\prime},r^{\prime}]}$.
@@
The lazy, eager, and our version of a relaxed approach are depicted in Figure 1 under the form of tilings of the contribution space. Note that the lazy and eager approaches can still be thought of as grouping contribution accounting, but do so along thin strips: the contribution of a long range of $y$’s to one element of $z$ in the lazy approach and that of a singe element of $y$ to a long range of $z$’s in the eager approach. The key to our relaxed tiling is using more balanced tiles since, as noted, the cost of a tile is given by the larger side (which is very expensive for thin tiles). The tiles still have to respect $r\leq l^{\prime}$ and one needs to ensure that the whole line of contribution to some $z_{t}$ has been dealt with by the time we return the value of $z_{t}$ - this is indeed the case for our tiling.
@@
2:for $i\leftarrow 1$ to L1L-1 do
3:  $U\leftarrow$ maximum power of $2$ that divides $i$ # the side of the $i$th gray tile
4:  $z_{i}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=y_{i}*\rho_{0}$ # finalize $z_{i}$ by accounting for newly available $y_{i}$ - red cell
5:  # account for the contribution of $y_{[i-U+1,i]}$ to $z_{[i+1,i+U]}$ - gray tile
6:  $z_{[i+1,i+U]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=\tau(y,[i-U+1,i],\rho,[i+1,i+U])$ # $\tau$ is defined as in Eq 3
7:  return $z_{i}$ # $z_{i}$ has been computed
8:  $unlock(y_{i+1})$ # $y_{i+1}$ becomes available
@@
[…]
11: return $z_{L}$ # $z_{L}$ has been computed
While the lazy and eager algorithms have been described, one needs to make explicit the order of processing the tiles in the relaxed setting since we still cannot use $y_{t}$ before $z_{t-1}$ has been computed. This is shown in Algorithm 1: return marks a value of $z$ having been computed and unlock represents a value of $y$ becoming available. At the beginning of iteration $i$ it always holds that $z_{i}=\sum_{j=1}^{i-1}{y_{j}\cdot\rho_{i-j}}$. We start the iteration by completing this with the freshly unlocked $y_{i}\cdot\rho_{0}$ at line 4 - this corresponds to the red tiles of Figure 1. Following this step $z_{i}$ is ready to be returned. Then, we account for the gray tile that has just been unlocked (line 6) - this tile corresponds to the contribution of $y_{[i-U+1,i]}$ to $z_{[i+1,i+U]}$ where $U$ is the side of the tile, namely the largest power of $2$ dividing $i$. Finally we return $z_{i}$ as being in its final form and get access to $y_{i+1}$ - we could have done this before dealing with the gray tile as well. Note that the relative ordering of the red and gray lines does not matter within an iteration, but it is important to finish up the first red and gray tiles before second ones, the second ones before the third ones and so on.